OpenStudy (anonymous):

Which shows the expression below in simplified form? (2 × 107) ÷ (5 × 104)

OpenStudy (anonymous):

@StudyTheWorld

OpenStudy (anonymous):

107 260 (Decimal: 0.411538)

OpenStudy (studytheworld):

2*107= 214 5*104=520 214/520 = 107/260 Answer: 107/260

OpenStudy (anonymous):

107/260*

OpenStudy (studytheworld):

Yes, and it's decimal form is 0.411538

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok thank you